emmavictoria:

Can someone help me with these 5 math questions?

Tyrion:

post em

emmavictoria:

Tyrion:

oof hold up i can recommend someone give me a sec

emmavictoria:

ok thanks!

Tyrion:

@Hero @Vocaloid

Tyrion:

@Shadow @Ultrilliam

Vocaloid:

in general, a number in scientific notation will take the form a * x^(n) where \[1 \le a < 10\] and n is a positive integer so, 1.0 * 10^(2) would count as scientific notation, but 11 * 10^2 would not. (11 * 10^2 would need to be re-written as 1.1 * 10^3)

Vocaloid:

so, let's look at your first problem: T = \[2\pi \sqrt{\frac{ 4.5*10^{-2} }{ 2.0*10^{3} }}\] start by simplifying the expression inside the radical, using your exponent rules then apply the radical by taking the square root of your previous result finally, since you have a 2pi on the outside, multiply the previous result by 2pi let me know what you get from that, or if you're still confused by any steps

emmavictoria:

I got 0.0298 or 2.98* 10^-2

Vocaloid:

good, that's what I got too (doesn't say anything about sig figs or rounding so I guess we can just round to 2 digits like you did)

Vocaloid:

for the next one, you can repeat the same process (multiply the numbers out, then convert to scientific notation by moving decimal places as necessary). let me know what you get or if you're getting stuck

emmavictoria:

Could you help me with the units?

Vocaloid:

oh yeah, sure notice how kg and kg cancel out so we're left with \[\frac{ 1 }{ \frac{ 1 }{ s^{2} } }\] if we re-write this fraction as a multiplication problem (change the division to a multiplication sign, then flip the second fraction), we get \[\frac{ 1 }~{ \frac{ s^{2} }{1 } }\] or just s^2

Vocaloid:

however, since that's inside the square root ---> take the square root to get s

emmavictoria:

Ok thanks! I got it now. For #2, would the ^2 be a part of the m/s?

emmavictoria:

Without the units, for #2 I got 6.96*10^6

Vocaloid:

yes, I got that value too for the units, yes, the ^2 gets applied to the m/s therefore it becomes ---> m^2 / s^2 with kilograms combined, it becomes (kg m^2) / s^2

emmavictoria:

Ok so does that mean the units for #2 are (kg m^2) / s^2?

Vocaloid:

yes

emmavictoria:

Ok, I'm going to try question 3

emmavictoria:

I got to (9x10^-9) * (9.6*10^-17)N, but I don't think that's right :/

Vocaloid:

you're on the right track, however, remember that it's (0.32)^2 not just 0.32, so you'll need to divide by 0.32 again your units are right (N)

emmavictoria:

Ohh, okay, that makes sense

emmavictoria:

3*10^-16?

Vocaloid:

hm, I got something a bit different this part was right (9x10^-9) * (9.6*10^-17), you just need to divide this by 0.32 ---> (9x10^-9) * (9.6*10^-17) / 0.32 = ?

emmavictoria:

2.7*10^-24?

Vocaloid:

yup, so 2.7 * 10^(-24) N is your solution

emmavictoria:

Alright thanks!

emmavictoria:

I'll try 4 and 5 on my own and let you know if I need any help! But thank you so much!

Vocaloid:

sure, just lmk if you're having trouble